# Modeling Human Behavior in a Strategic Network Game with Complex Group Dynamics

### v2

@@ C.4 Resampling
The last step of the EPDM algorithm (Algorithm 1 from the main paper) consists of resampling the parameterizations evaluated in the last generation $\Pi(G-1)$ based on fitness. This is done by randomly selecting $N$ parameterizations from the list $\Pi(G-1)$. In selecting each parameterization, the probability that parameterization $\pi$ is chosen is given by $\frac{|\Gamma^{\pi}(g)|}{\sum_{\pi^{\prime}}|\Gamma^{\pi^{\prime}}(g)|}$. Note that a parameterization can be chosen more than once.
Appendix D Error Functions
Both the PSO and EPDM algorithms for learning parameterizations from data require an error or distance function (e.g., Eq. 1 in the main paper). This function takes as input two token allocation profiles, and returns a value specifying the difference between them. We experimented with several such functions, including mean-squared error (MSE) and a custom built function, called property scoring (PS). In PS, the error is based on how well the two token allocation profiles match across a variety of different properties. More specifically, the PS error function scores pairs of token allocation profiles using a set of properties. The distance between the actual allocations and the proposed allocations is decreased for keeping the right number of tokens, giving the right number of tokens in total, giving tokens to the right players, giving the right number of tokens to the right players, taking at the right times, taking the right number of tokens, and taking from the right player. The distance function is increased for proposing to give to a player from which tokens were actually taken or taking tokens from a player for which tokens were actually given.
Exact specifications of the property scoring distance function are given in the file EPDM.cpp (see the function scoreProposedAllocation) of the supplied code.
Because the PS error function always performed better than MSE, we used this scoring function in the training of all agent models.
Appendix E Metrics
In this section, we define how we computed each of the metrics used in the paper to evaluate population dynamics. These metrics are derived from the known quantities of the JHG, including the transaction matrix $X$, the popularity vector ${\mathcal{P}}$, and the influence matrix $\mathcal{I}$. Additionally, to capture longer-term (non-discounted) interactions between players, we use the impact matrix. We define the impact matrix prior to discuss each of the metrics used in the paper.
We define the economic actions taken by a player at time $\tau$ as a matrix $X(\tau)$. Some metrics, discussed below, use both positive and negative connections while others only one or the other. To distinguish between each case while trying to simplify explanations we distinguish between $X(\tau)$, $X^{+}(\tau)$, and $X^{-}(\tau)$ as all connections, positive connections, and negative connections. We apply this to the different derived matrices as well, such as the influence matrix or impact matrix presented below.
E.1 The impact matrix
To examine the evolution of friendships and interaction patterns throughout the Junior High Game, we construct the impact graph, a network representation that aggregates player interactions over some duration of the game. This graph encodes social mixing patterns by weighting interactions based on their frequency, the relative popularity of the involved players at the time of interaction, and the corresponding economic action coefficient.
To calculate the impact matrix, we first calculate $J^{\prime}_{ij}(\tau)$ representing the interaction strength between players $i$ and $j$ at time step $\tau$. It is defined as:
$J_{ij}(\tau)=\begin{cases}\mathcal{P}(\tau)X_{ij}(\tau)c^{keep},&\text{if }i=j\\ \mathcal{P}(\tau)X_{ij}(\tau)c^{give},&\text{else if }x_{ij}\geq 0\\ \mathcal{P}(\tau)X_{ij}(\tau)c^{take},&\text{else if }x_{ij}<0\end{cases}$ (2)
where $\mathcal{P}_{i}(\tau)$ represents the popularity of player $i$ at time $\tau$ adjusting the weight of interactions based on social influence. $X_{ij}(\tau)$ is the economic action between players $i$ and $j$ with $c^{give}$, $c^{keep}$, and $c^{take}$ representing the economic action coefficients that assign weights based on the type of interaction (giving, keeping, or taking) respectively.
The aggregated impact matrix over a time interval $[q,r]$ is given by:
$J^{\prime}(q,r)=\sum_{\tau=q}^{r}J(\tau)$ (3)
which sums the impact matrices over all time steps within the interval. We then normalize the interaction weights so that we can compare relative influences across different context as follows:
$J^{\prime}(q,r)=\frac{J(q,r)}{\|J(q,r)\|_{1}}$ (4)
where $\|J^{\prime}(q,r)\|_{1}$ is the L1 norm of the matrix, ensuring that the sum of the magnitude of all entries in $J(q,r)$ is constrained within a unit scale.
Additionally, we often look at the impact of just one round. That is, if $q=r$, we use the notation $J^{\prime}(r)$ to mean $J^{\prime}(q,r)$ to simplify understanding of our equations.
E.2 Metrics for wealth and power
Because the JHG uses popularity (an approximation of Katz centrality [24]) as the primary commodity, we utilize the mean popularity, $\mu_{\mathcal{P}}$, as a metric to quantify the productivity of a society. To apply this measure, we first constructed a popularity vector $\mathcal{P}=[\mathcal{P}_{0},\mathcal{P}_{1},\ldots,\mathcal{P}_{n-1}]$ representing the popularity of each agent $i$ at a particular time $\tau$. We then define average popularity as follows:
$\mu_{\mathcal{P}}(\tau)=\frac{1}{n}\sum_{i=1}^{n}\mathcal{P}_{i}(\tau)$ (5)
We additionally employ the Gini Index, $G(\tau)$, as a quantitative metric to assess inequality of popularity across a society. The Gini Index, originally developed as a measure of income inequality [16], ranges from 0 to 1, where 0 indicates perfect equality (i.e., all individuals possess equal popularity) and 1 denotes maximal inequality (i.e., one individual possesses all of the popularity while others have none). $G(\tau)$ was then computed using the following standard formula:
$G(\tau)=\frac{1}{2n^{2}\mu_{\mathcal{P}}(\tau)}\sum_{i,j=0}^{n-1}|\mathcal{P}_{i}(\tau)-\mathcal{P}_{j}(\tau)|$ (6)
where $n$ is the number of agents and $\mu_{\mathcal{P}}$ is the mean of the values in vector $\mathcal{P}$. In our analysis, we used the Gini Index (as opposed to the variance of the population) because it is a preferred measure used in measuring inequality in societies in economics [38, 2] and sociology [28].
We used average popularity and the Gini Index as an interpretable and sensitive measures for assessing the mean and spread of popularity in a society.
E.3 Metrics for quantifying of economic behavior
To help understand how individuals in a society interact with each other, we analyze their economic actions using the percent of give, keep and take allocations over societies of each type of agent. Precise definitions of these metrics are as follows:
[…]
where $X_{ij}(\tau)$ represents the token allocations of player $i$ toward player $j$ at time $\tau$. The diagonal of the transaction matrix $X(\tau)$ specifies agent keeping, while positive non-diagonal values represent giving and negative non-diagonal values represent taking. $T$ is the number of rounds in the game, $|I|$ is the cardinality of the set $I$ (specifying the number of players in the game), and $\delta$ is the Kronecker delta function.
E.3.1 The evolution coefficient
To capture the evolution of economic behaviors in the JHG over time, we use the evolution coefficient. The evolution coefficient quantifies structural changes in the impact matrix over time. Given that the impact matrix $J^{\prime}(q,r)$ represents normalized interaction patterns over a specified window. This coefficient, denoted $C_{d}$, is computed by evaluating differences in consecutive impact matrices, shifted by a delay $d$:
$C_{d}=\sum_{\tau=0}^{r-p-1}\sum_{ij}\left|J^{\prime}_{ij}(\tau)-J^{\prime}_{ij}(\tau+d)\right|$ (10)
In words, the evolution coefficient quantifies the cumulative absolute differences between the impact matrices over a specified number of rounds $d$. A higher value of $C_{d}$ indicates more substantial shifts in the interaction structure, suggesting dynamic social mixing or evolving economic strategies among players. Conversely, lower values suggest stability in social interactions and player behaviors.
We note that, in the computation of the Mahalanobis distance, the evolution coefficient is measured as the average change in token allocations across one and two rounds of play.
E.4 Mixing patterns: Measuring reciprocation
Reciprocity is an important component of cooperative dynamics within agent societies [31]. High reciprocity suggests that players consistently return favors and maintain balanced exchanges, whereas low reciprocity indicates asymmetric relationships, where some players give or take more than they receive. We define the pairwise global reciprocity between agents $i$ and $j$ over the course of an interaction as:
$R^{global}_{ij}=\begin{cases}0&\text{if }i=j;J_{ij}(r_{0},r_{f})\leq 0;J_{ji}(r_{0},r_{f})\leq 0\\ J^{\prime}_{ji}(r_{0},r_{f})\frac{J_{ij}(r_{0},r_{f})}{J_{ji}(r_{0},r_{f})}&\text{else if }J_{ij}(r_{0},r_{f})<J_{ji}(r_{0},r_{f})\\ J^{\prime}_{ji}(r_{0},r_{f})&\text{else if }J_{ij}(r_{0},r_{f})\geq J_{ji}(r_{0},r_{f})\\ \end{cases}$ (11)
Here, $J_{ij}(r_{0},r_{f})$ denotes the cumulative reciprocated impact from agent $j$ to agent $i$ over the interval $[r_{0},r_{f}]$, where $r_{0}$ if the first round of the game and $r_{f}$ is the final round of the game. The metric $R^{global}_{ij}$ thus captures the extent to which agent $j$ reciprocates the contributions made by agent $i$, normalized by the directionally lesser interaction.
We use the overall global reciprocity coefficient to define overall reciprocity. It is summarizes the aggregate reciprocity across all agent pairs in the population as follows:
$R^{global}_{coeff}=\frac{\sum_{ij}{R^{global}_{ij}}}{n(n-1)}$ (12)
where $n=|I|$ denotes the number of players in the game. This coefficient provides a population-level estimate of the prevalence of mutual exchanges, weighted by the magnitude of impact transferred between agents over the duration of the game.
In addition to this cumulative measure, we also define immediate reciprocity to capture short-term reciprocation dynamics. The pairwise local reciprocity at round $r_{m}$ is defined as:
$R^{local}_{ij}=\begin{cases}0&\text{if }i=j;J_{ij}(r_{m})\leq 0;J_{ji}(r_{m})\leq 0\\ J^{\prime}_{ji}(r_{m-1})\frac{J_{ij}(r_{m})}{J_{ji}(r_{m-1})}&\text{else if }J_{ij}(r_{m})<J_{ji}(r_{m+1})\\ J^{\prime}_{ji}(r_{m-1})&\text{else if }J_{ij}(r_{m})\geq J_{ji}(r_{m-1})\\ \end{cases}$ (13)
The corresponding local reciprocity coefficient is:
$R^{local}_{coeff}=\frac{\sum_{ij}{R^{local}_{ij}}}{n(n-1)}$ (14)
This formulation evaluates how frequently short-term exchanges are reciprocated from one round to the next, capturing the temporal responsiveness of agents to recent interactions.
E.5 Mixing patterns: Measuring connectivity using density and entropy
Graph density provides a measure of how interconnected a network is at a given time, relative to the maximum possible number of directed edges (excluding self-loops). Just to note, when average across the whole network average in/out degree are proportional to density and therefore we exclude the redundant comparison. Density is defined as:
$D(\tau)=\frac{1}{n(n-1)}\sum_{i,j=0;i\neq j}^{n-1}A^{+}_{ij}(\tau)$ (15)
Here, $A^{+}_{ij}(\tau)$ represents an adjacency matrix which is $1$ if a positive economic action from node $i$ to node $j$ took place in round $\tau$. Otherwise it is 00. The denominator $n(n-1)$ reflects the maximum number of actions in a loop-free graph.
While density describes the average number of associates a player is interacting with, entropy quantifies the distribution of a player’s actions among agents. We define a normalized entropy score that captures how evenly distributed its outgoing edge weights are. For node $i$, let the set of positive outgoing edge weights be $\{X^{+}_{ij}\}_{j=1}^{n}$. Define the normalized edge weight distribution as:
$w_{ij}=\frac{X_{ij}^{+}}{\sum_{k=1}^{n}X_{ik}^{+}}$ (16)
The entropy for node $i$ is then given by:
$H_{i}=-\sum^{n-1}_{j=0;X^{+}_{ij}>0}w_{ij}\log w_{ij}$ (17)
The maximum possible entropy $H_{i}^{\text{max}}$ occurs when all outgoing edges from node $i$ are equally weighted, i.e.,
$H_{i}^{\text{max}}=\log k_{i},$ (18)
where $k_{i}$ is the out-degree of node $i$ (number of positive-weight outgoing edges). Then entropy is then computed as:
$H_{i}^{\text{norm}}=\begin{cases}0&\text{if }\sum_{j}X^{+}_{ij}=0\text{ or }k_{i}=1\\ \frac{H_{i}}{H_{i}^{\text{max}}}&\text{otherwise}\end{cases}$ (19)
Finally, the average entropy across all nodes provides a summary statistic for the overall spread of edge weight distributions in the network:
$H^{\text{avg}}=\frac{1}{n}\sum_{i=1}^{n}H_{i}^{\text{norm}}$ (20)
E.6 Measuring group formation: Polarization
Finally, we desire to quantify how players cluster and separate. Known metrics for (partially) quantifying such behaviors include the clustering coefficient [47, 29] and modularity [30]. However, many metrics either (1) are not designed for signed, weighted, and time-varying graphs, (2) do not provide comparisons of grouping behavior from graph to graph (e.g., modularity measures how well a community partition describes a graph, not how divided or grouped a graph is), or (3) conflate multiple concepts (e.g., high density tends to produce a high clustering coefficient and thus may not indicate how members of society actually group and segregate). Thus, in this paper, we use a customized metric called polarization to quantify how individuals separate into different groups and sub-groups. Related somewhat to the idea of modularity, polarization measures how much individuals would consistently prefer individuals within a designated group as opposed to individuals outside of that group.
Polarization measures the maximum preference separation between two distinct (and sufficiently important) sub-groups in the society. Preference separation is a method for measuring how separated two distinct sub-groups within society are in a weighted and signed graph. Specifically, it measures how much individuals in a sub-group would prefer (based on immediate connections) any one member of their own sub-group to any one member of the other sub-group. Preference separation between two groups is highest when (1) nodes with each sub-group are well connected, (2) connections between the two sub-groups are weaker than connections within the sub-groups, and (3) the sub-groups cover a larger proportion of the entire society.
Before giving details for how preference separation and polarization are computed, we first define several quantities. Let
$m_{i}=\sum_{j}|\mathcal{I}_{ji}|$ (21)
be the magnitude of influence (both positive and negative) flowing to individual $i$ from all other individuals in society. Furthermore, let
$w_{i}=\frac{m_{i}}{\sum_{j\in I}m_{j}},$ (22)
where $I$ is the set of all individuals in the society, be the total proportion of influence that is flowing to individual $i$.
Next we define a set of dueling sub-groups in society by comparing how society would partition itself if all individuals had to choose between individuals $i$ and $j$ (where $i\neq j$. Let
$G_{ij}=\{i\}\cup\{k\in I:\mathcal{I}_{ik}>\mathcal{I}_{jk}+\delta\},$ (23)
where $\delta=0.04m_{k}$, denote the group of individuals in society that would prefer $i$ over $j$. Similarly, $G_{ji}$ denotes the group of individuals that prefer $j$ over $i$. Finally, let $G_{ij}^{\sim}=I-(G_{ij}\cup G_{ji})$ denote the set of individuals in society that are indifferent between individuals $i$ and $j$. Finally, let $W_{ij}=\sum_{k\in G_{ij}}w_{i}$ denote the collective weight of group $G_{ij}$.
Preference separation between sub-groups $G_{ij}$ and $G_{ji}$, denoted $\mathcal{S}(G_{ij},G_{ji})$, measures the proportion of individual in $G_{ij}$ and $G_{ji}$ that would prefer to stay in the same group if the choice were between other individuals from the respective groups (other than just $i$ and $j$). Formally, it is computed as follows:
$\mathcal{S}(G_{ij},G_{ji})=\left(\frac{\sum_{k\in G_{ij}}\sum_{\ell\in G_{ji}}|G_{ij}\cap G_{k\ell}|+|G_{ji}\cap G_{\ell k}|+0.5|G_{k\ell}^{\sim}|}{|G_{ij}||G_{ji}||I|}\right)\left(\frac{{\rm min}(W_{ij},W_{ji},0.2}{0.2}\right),$ (24)
where ||$\cdot$|| ⋅ | denotes the cardinality of a set. Note that the last item in the numerator of left term gives half credit when individuals are neutral between the two groups for a particular pairing. The right component of the equation scales down preference separation if one of the groups is too small (determined as having a weight of less than 0.2).
Note that measures of separation tend to be between 0.5 and 1. As such, we base polarization on a normalized version of preferences separation, given by $\mathcal{S}^{\prime}(G_{ij},G_{ji})=\frac{\mathcal{S}(G_{ij},G_{ji})-0.5}{0.5}$.
Once preference separation is computed for all pairs of sub-groups, polarization is determined to be the max preference separation for pairs of subgroups. That is, polarization (denoted $\Gamma$) is given by $\Gamma=\max_{i,j}\mathcal{S}^{\prime}(G_{ij},G_{ji})$. The constraint to only consider subgroups of a particular size ensures that the preference separation between the groups could have a meaningful impact on society.
All of the above description omit the time $t$ – polarization is computed for each time period. The actual metric we report is the average polarization over all rounds considered. We note that this metric is not computationally efficient in this formulation, but it can be computed and utilized effectively for small societies such as those we consider in this paper.
E.7 Summary metric: Mahalanobis distance
Each of the above metrics represent important aspects of population dynamics. To further compare and contrast human and agent populations, we aggregate these metrics into a multivariate Gaussian. That is, we quantify the dissimilarity between populations with respect to these metrics using the Mahalanobis distance [27]. Unlike the Euclidean distance, which assumes independence across dimensions, the Mahalanobis distance incorporates the covariance structure of the data, making it well-suited for comparing complex, interdependent societal metrics.
Given two societal types, we compute the Mahalanobis distance between them as:
$D_{M}(\mathbf{x})=\sqrt{(\mathbf{x}-\boldsymbol{\mu})^{T}\Sigma^{-1}(\mathbf{x}-\boldsymbol{\mu})}$ (25)
where $\mathbf{x}$ is the vector of observed societal metrics for a given type, $\boldsymbol{\mu}$ is the mean vector of another societal type, and $\Sigma$ is the covariance matrix of the distribution of human behavior.
To complement the Mahalanobis distance, we compute the Chi-Squared p-value, which indicates the statistical significance of the observed differences. Since the squared Mahalanobis distance follows a Chi-Squared distribution with degrees of freedom equal to the number of dimensions, the p-value provides a probabilistic interpretation of the similarity between societal types. Lower p-values suggest significant differences between the compared distributions, while higher p-values indicate that the societal structures are statistically indistinguishable.
This methodology enables us to identify when two societies exhibit fundamentally different organization patterns and to characterize the extent of these differences.
Appendix F Experiment A
Additional implementation details and results from the first experiment reported in the main paper:
  • We use the same JHG parameter settings as was used by Skaggs et al. [41]. That is, $\alpha=0.2$, $\beta=0.5$, $c_{\rm give}=1.3$, $c_{\rm keep}=0.95$, $c_{\rm take}=1.6$. Furthermore, each agent had $2*|I|$ tokens to allocate in each round.
  • eCABs were evolved with a single set of genes (rather than the 3 sets used by Skaggs et al. [41]) and were trained for 200 generations with variable initial probabilities. Simulation results were conducted with randomly selected parameterizations from the 100 parameterizations that were in the final generation.
  • Additional precision along with standard for average economic actions in each society are provided in Table 4
  • No GPU was used in the study. No high-powered computational equipment were used (only standard computing equipment with multiple cores). Training using the PSO algorithm was done on a 32-core machine, which took about 3 hours per run. Training using the EPDM algorithm took about 30 minutes on a machine with 16 cores.
Agent-Society % Give % Keep % Take
@@ Appendix F Experiment A
Table 4: Mean percent of each economic actions in human and agent societies. $\pm$ values show show standard error of the mean.
Appendix G Experiment B: User study details
Additional information about the second experiment (user study) reported in the main paper:
  • The user study was approved by the authors’ institutional review board (it was given exempt status). Participants gave consent to participated in the study through the approved process.
  • Participants were paid approximately \$16 (USD) per hour while participating in the study. A high-score list (based on final popularity in each game) was kept to motivate participants to perform well.
  • Games were played using the publicly available online platform published by Skaggs et al. [41] (permission was obtained from the authors of that work to use the platform for this study). This online platform can be found at juniorhighgame.com.

### v3

@@ 3.2. Learning Parameterizations
After all generations are completed, the last step of the EPDM algorithm (Algorithm 1) consists of resampling the parameterizations evaluated in the last generation $\Pi(G-1)$ based on fitness. This is done by randomly selecting $N$ parameterizations from the list $\Pi(G-1)$. In selecting each parameterization, the probability that parameterization $\pi$ is chosen is given by $\frac{|\Gamma^{\pi}(g)|}{\sum_{\pi^{\prime}}|\Gamma^{\pi^{\prime}}(g)|}$. Note that a parameterization can be chosen more than once.
Table 1. Metrics used to compare the population dynamics of human and agent societies.
Category Description Metrics
Economic Behavior High-level actions (give, take, keep) taken by individuals in society % Give, % Take, % Keep, Evolution Coefficient
Mixing Patterns Connections across society, including dyadic, triadic, and group relationships Reciprocity, Density, Entropy, Polarization
4. Evaluating population dynamics
Our goal is to model human behavior in the JHG so as to satisfy two criteria. First, the population dynamics of agent collectives should mirror the population dynamics of human societies. Second, individual agents should, ideally, also exhibit human-like behavior. In this section, we evaluate four models, created by combining parameter-based models (hTFT and hCAB111hCAB and hTFT denote parameterizations of CAB and TFT models, respectively, learned from human data.) with modeling algorithms (PSO and EPDM), with respect to the first criteria. In Section 5, we address the second criteria.
We note that human behavior is likely impacted by group size. Understanding human behavior given all group sizes is valuable. However, given that our data sets (Section 4.2) consist of small-scale human groups (5-12 individuals), we only consider groups of this size in the evaluations made in this paper. While we anticipate that human behavior in larger groups could be modeled with the same algorithmic mechanisms, we leave such evaluations to future work.
4.1. Experiment Design
We simulate societies of hTFT-PSO, hTFT-EPDM, hCAB-PSO, and hCAB-EPDM agents in the JHG, and compare their dynamics to those of games played by human players. We also consider two baseline agents: (1) Random, which randomly allocates tokens with the percentages of give, take, and keep that match the percentages observed in the training set (Section 4.2) and (2) eCAB agents (skaggs2024fostering). eCABs are CAB agents with parameterizations learned via evolutionary simulations rather than from human data. eCABs display somewhat effective behavior in the JHG, though their strategies are distinct from those used by humans skaggs2024fostering.
For simulations with hTFT-EPDM, hCAB-EPDM, and eCAB agents, parameterizations were randomly selected from the learned pool of strategies for each game. On the other hand, for hTFT-PSO and hCAB-PSO, the parameterizations to which the PSO algorithm converged in four separate runs of the PSO algorithm were used.
4.2. Data Sets
4.3. Metrics to Evaluate Population Dynamics
We evaluate population dynamics using the metrics summarized in Table 1. These metrics are grouped into four categories, each designed to evaluate a particular aspect of population dynamics. Metrics in the first category quantify the wealth and power wielded by society members. We use Mean Popularity, an approximation of Katz centrality (katz1953new), and Gini Index, which measures the distribution of wealth across society. Economic behavior summarizes player actions across society, including the percentages of token allocations of each type (keep, take, and give) and the average change in player token allocations from round to round (Evolution Coefficient). Next, we assess societal mixing patterns, or the way that players both join together in pairs and groups, and fragment and segregate. These mixing patterns are quantified using two measures of reciprocity, density, entropy, and polarization. Finally, we summarize the dissimilarity between agent and human populations across all these metrics using Mahalanobis distance (mahalanobis1936generalized). Formal definitions of each metric, as well as motivation for why we chose these (as opposed to other) metrics, are provided in SM-5.
4.4. Results
Figure 2 displays measures of human and agent population dynamics with respect to individual metrics. We discuss each by category.
(a)                  (b)                  (c)
Society % Give % Keep % Take
@@ 4.4. Results
[…]
Figure 2. Comparisons of human and agent societies. Dots (a-c), numbers (d), and lines (e-f) denote the mean, error ovals (which do not account for covariance), $\pm$, and error ribbons show the standard error.
Wealth and Power: Figure 2a compares agent and human populations with respect to Mean Popularity and Gini Index. In these two regards, hCAB-EPDM populations most closely mirror those of humans. Random populations tend to have both low Gini Index and low mean popularity. hTFT-EPDM populations tend to have slightly higher Gini Index and lower popularity, while hTFT-PSO, hCAB-PSO, and eCAB populations all have higher Mean Popularity and lower Gini Index. While high Mean Popularity and lower Gini Index is desirable (though this does not necessarily indicate robust behavior in general), our objective is to identify models that mimic human behavior. With respect to Mean Popularity and Gini Index, the CAB model coupled with EPDM parameter modeling results in the model that most closely reflects human population dynamics.
Economic Behavior: In the JHG, wealth outcomes are the result of economic behavior. eCAB agents and algorithms that model parameters using PSO all have higher Mean Popularity and lower Gini Index than human populations. These agents allocate most of their tokens to giving. They keep and take little, if any (Figure 2d). On the other hand, humans and agents modeled with EPDM have higher variability in strategies across players and tend to keep and take more tokens. This produces greater societal inequality and less overall prosperity. As such, agents produced using EPDM have token allocation distributions that closer to those of humans.
Mixing Patterns: We measure two forms of reciprocity: Immediate Reciprocity (reciprocation from one round to the next) and Overall Reciprocity (reciprocation over a complete game). In the JHG, humans have moderate levels of both forms of reciprocity (Figure 2b). EPDM agents exhibit similar levels of reciprocity, with hCAB-EPDM’s rates being only slightly higher than people’s. On the other hand, eCABs and PSO reciprocation rates are far higher than those of the humans in evaluated in our test set.
The density and entropy measures, shown in Figure 2c, provide information about the number (as measured by Density) and strength (as measured by Entropy–this metric measures whether all trading partners are valued similarly) of trading partners. In these regards, hCAB-EPDM agents once again tend to be very similar to those of human players. hCAB-PSO agents have similar Density, but have higher Entropy. On the other hand, agents modeled using the TFT model tend to be less similar to humans in both regards. Thus, hCAB agents mirror these human behaviors better than TFT agents, with hCAB-EPDM being the closest.
Finally, we consider Polarization (Figure 2f), which describes the degree to which players segregate into groups. Since the CAB algorithm is based on group formation, while TFT is not, it is not surprising to observe that CAB agents display higher Polarization than TFT agents. These higher levels are similar to the Polarization measures we observe in human games in the test set. This again shows that the CAB model can be tuned to better model human dynamics than the TFT (matching) model.
Table 2. Mahalanobis distance (lower is better) between human and agent populations over all metrics. $p>0.05$ indicates that we cannot reject the null hypothesis (that they come from the same distribution).
Algorithm Mahalanobis Distance Chi-Square Test
Random 18.822 $p<0.001$
eCAB 7.089 $p<0.001$
hTFT-PSO 11.565 $p<0.001$
hTFT-EPDM 7.656 $p<0.001$
hCAB-POS 9.216 $p<0.001$
hCAB-EPDM 3.393 $\mathbf{p=0.486}$
Summary Metric: Table 2 gives the Mahalanobis distance between the various agent populations and our human population across the eleven individual metrics. Of the six agent populations, hCAB-EPDM populations most closely mirror human societies from our test set. Indeed, a Chi-Square test comparing human and hCAB-EPDM populations produces a p-value ($p=0.486$) indicating that we cannot reject the null hypothesis (that they come from the same distribution), thus indicating a somewhat close fit across these measures. All other agent populations were statistically distinct from human populations ($p<0.001$ in all cases).
These results show that the CAB algorithm paired with EPDM parameterization modeling produces agent populations whose dynamics are surprisingly close (with some variation–notably the Evolution Coefficient) to those of human populations in the JHG (first criteria). Is this relatively similar behavior limited to population dynamics, or does it also carry over into individual behaviors? In the next section, we complement the quantitative comparisons of human and agent populations with a user study designed to evaluate whether individual hCAB-EPDM agents plausibly exhibit human behavior (second criteria).
5. Interacting with human players
To further illustrate the behavior of hCAB-EPDM agents (hereafter referred to as hCABs), we conducted a user study in which experienced human players interacted with hCABs in the JHG. From this study, we (1) illustrate the behavior of individual hCABs when interacting with humans and (2) quantify how well human participants were able to distinguish hCABs from humans.
5.1. User Study Design
We recruited eight people, each of which had prior experience playing the JHG, to participate in the study. The study proceeded in a series of four periods. In each period, the participants were randomly divided in half. Each group of four participants was paired with four (randomly selected) hCABs to form two games played by eight players each. Thus, eight games were played across the four periods. Games lasted for 21, 20, 17, and 15 rounds in Periods 1-4, respectively (players were not told game lengths). For consistency, only the first 15 rounds of each game were used in our data analysis.
5.2. Results
Figure 3. Mean popularity over time of humans and hCABs in the user study. Error ribbons show the standard error.
[…]
A comparison of the mean popularity over time of humans and bots in the user study.
Table 3. The percentage of tokens allocated to giving, keeping, and taking by humans and hCABs in the user study.
Society % Give % Keep % Take
Humans 60 25 15
hCABs 75 20 5
Table 4. The percentage of players correctly identified as human or bot by human participants verses random chance.
Guesser % Correct Designation
Humans 56.7
Chance 57.1